### v3

@@ 6.2 Counterfactual selection policies for mentors
Applicants selected by mentors and randomized to be the control group had a significantly higher likelihood of securing a tech job compared to those who were not selected (see Figure 2). This suggests that mentors tend to prefer highly promising applicants who are likely to secure a tech job irrespective of their participation in Mentoring. In this section, we examine alternative selection strategies that mentors could adopt and their potential implications for the effectiveness of the mentoring program. Specifically, we consider four scenarios: (i) Highest CATE – targeting applicants predicted to have the highest treatment effects, (ii) Most Promising – targeting applicants expected to have the highest outcomes without participating in Mentoring, (iii) Least Promising – selecting applicants expected to have the lowest outcomes if they do not participate, and (iv) Highest Chances if Treated – where mentors select participants that have the highest predicted outcomes if they participate in Mentoring; additionally, we include a case in which mentors select at random (Random Choice) and the observed allocation (Observed) as benchmarks. In each policy, we assume that 36% of applicants participate in Mentoring, which corresponds to the share of applicants that were selected by mentors. [^17] We consider a higher number of participants than the current number of spots in Mentoring to amplify the impact of policies on outcomes across all applicants; the ordering of policies does not depend on the share of applicants admitted to Mentoring as long as this number is constant across policies.
We follow the same off-policy approach as previously. We consider the dataset of all applicants to Mentoring, which consists of 151 individuals who participated in the program, 147 mentors selected but randomized to the control group, and 468 non-selected applicants. We split data into train and test with equal probability. We use the training dataset to estimate the mapping from applicants’ characteristics to assignments. To do that, we train the Cox proportional hazard model in the training data and predict outcomes with and without treatment in the testing data. Then, we use these predictions to determine which applicants get selected under each policy. To evaluate policies, we compute AIPW scores analogously to the optimal policy evaluation (Appendix J.1): using only testing data, we train the propensity model (logit model) and a new outcome model (Cox hazard model) using cross-fitting. We compute 100 cross-fitting rounds and take the median outcome. Table 6 shows the results.
Table 6: Mentors’ selection policies.
Policy Mean ATE Baseline outcome treated group Mean outcome all applicants
Highest CATE 0.178 (0.012) 0.187 (0.008) 0.232 (0.005)
Most Promising 0.062 (0.012) 0.285 (0.009) 0.192 (0.005)
Least Promising 0.127 (0.007) 0.055 (0.005) 0.211 (0.005)
Highest Chances if Treated 0.130 0.012 0.237 (0.008) 0.218 (0.005)
Observed 0.096 (0.012) 0.275 (0.008) 0.205 (0.005)
Random 0.094 (0.011) 0.191 (0.008) 0.203 (0.005)
Note: Outcomes and treatment effects across different mentors’ selection rules. Mean ATE is estimated using AIPW estimator with cross-fitting. Baseline outcome treated group is the predicted outcome without treatment for the group that receives treatment in the considered policy. Mean outcome all applicants corresponds to the average probability of getting a job in tech across all applicants, where the group assigned to treatment under considered policy receives treatment, and the remaining applicants do not participate. Standard errors clustered at the applicant level.
We find that prioritizing applicants with the highest anticipated treatment benefits results in higher average outcomes for all applicants and a more significant treatment effect than other strategies. Using the Highest CATE policy, 23% of applicants secure a tech job, a noticeable improvement over the 20% success rate under the Random or Observed policies. The average treatment effect under Highest CATE is an 18 percentage points increase in the probability of having a tech job for the treated group. Interestingly, targeting applicants with the highest chances of having a job if not treated (Most Promising) leads to treatment effects that are lower than selecting applicants randomly. Moreover, a policy aimed at the seemingly less promising applicants (Least Promising) — which goes against current mentor preferences — produces an average treatment effect of 13 percentage points. This results in an average outcome for all applicants surpassing those achieved by the Most Promising and Random policies. The policy of targetting applicants that have the highest chances of getting a tech job if treated (Highest Chances if Treated) results in a 13 percentage points average treatment effect, which is similar to Least Promising.
74% of applicants assigned to Mentoring under Highest CATE are also assigned to the program under Highest Chances if Treated, which is expected considering the fact that these two policies are most promising. However, there is a low overlap between Highest CATE and Least Promising only 7%. Furthermore, there is a substantial overlap between the observed selection by mentors and the Highest CATE policy – 54%, but a very low one between Observed and Least Promising, only 4%. These observations suggest that a group of applicants has low chances of securing a tech job when not treated and are systematically overlooked by mentors, even though their benefits from the program would be high. We analyze the overlaps and the characteristics of selected candidates in Appendix M.
6.3 Simple priority rules
Optimal assignment policy increases the average probability of getting a tech job across all applicants compared to the status quo. However, a non-parametric policy requiring constraint integer optimization is difficult to implement in Dare IT’s context. In this section, we examine whether a simple priority rule based on heuristics can enhance the average effectiveness of Dare IT’s programs.
Consider the admission policy wherein Dare IT first screens participants based on their priority level and then randomly assigns them to either the Mentoring or Challenges program (with the probabilities proportional to the number of spots in either of the programs). We propose the following priority groups: Priority 1 – Applicants who are from small towns, do not have a degree in a STEM field, and do not have a master’s or postgraduate degree; Priority 2 – Applicants who are from small towns, do not have a STEM degree, but have a master’s or postgraduate degree; Priority 3 – Applicants who are not from small towns and do not have a STEM degree; Priority 4 – Applicants who have a STEM degree.
The proposed priority groups are non-overlapping and together encompass all applicants. Thus, under this proposed admission policy, Dare IT would first consider the applicants in the first group. If there is more capacity after admitting the applicants in the first group, they will move to the second group, and so on. We are interested in evaluating whether the groups of applicants with high priority have higher predicted treatment effects than groups with lower priority levels and whether the high-priority groups are large enough to fill in all the existing spots in the two programs.
Given that this policy is defined using heuristics instead of data, we do not split the dataset for policy estimation and separate evaluation. Instead, we use the entire dataset to evaluate this pre-defined policy. To estimate the value of the policy, we consider both experimental populations and the remaining applicants. For each applicant, we estimate outcomes under three conditions: Mentoring, Challenges, and Out of Dare IT. We do that using cross-fitting as described in Appendix J.1. The outcome model is the proportional hazard Cox model, and the propensity model is the multinomial logit model; we combine them to obtain AIPW estimates. Admitted applicants are assigned with probability 45% to Mentoring and 55% to Challenges (reflecting the proportions of treated subjects in the dataset). Once we have the allocation, we compute the average treatment effect per group using the AIPW predictions. Figure 5 shows the results.
Figure 5: Evaluation of simple policies.
Note: Estimates of the average treatment effect per priority group; AIPW estimates obtained using cross-fitting. The treatment effect is the percentage point increase in the probability of a tech job. The red dotted line is the average treatment effect across the entire population. Whiskers around point estimates represent confidence intervals. Standard errors are clustered at the participant level.
We find that groups 1 and 2 have predicted average treatment effects that are higher than the average across the entire population of applicants (red scatter line) and higher than groups 3 and 4. Furthermore, group 1 encompasses 14% of applicants and group 2 10%. The combined capacity of Mentoring and Challenges is 28%. Thus, the two groups can almost fill in all the spots in the two programs. We find that this simple priority rule can increase Dare IT’s effectiveness.
@@ 7 Conclusion
We show that a mentoring program and an online training program called Challenges aimed at developing labor market signals increase the probability of having a job in the technology sector by more than 40% four months after the end of the program. Furthermore, the online training program developed during this research project is cheap to operate, at approximately \$15 per person, and suitable for scaling up, suggesting that it can have an impact extending beyond the current setting.
@@
Our findings reveal strong patterns of heterogeneity in treatment effects. The benefits derived from the two programs depend on a number of participants’ characteristics including the educational background, length of professional experience, type of residence, and age. For example, we find that older participants that come from small towns have particularly high treatment effects in both programs, especially in Mentoring. Furthermore, certain groups benefit distinctly from one program over the other. Specifically, participants with short professional experience benefit from Challenges, while those with over five years of experience from Mentoring.
@@
For the foreseeable future, the demand for the two programs, particularly for Mentoring will exceed the number of available spots, and Dare IT will grapple with the question of whom to admit. We propose an admission policy that maximizes Dare IT’s impact, by prioritizing applicants with the highest predicted treatment effects. Consistent with the high degree of heterogeneity in the effectiveness of the two programs we find that such targeted admission policies can substantially increase Dare IT’s effectiveness. Specifically, we estimate a seven percentage points increase in the average probability of getting a job in the technology sector across all applicants. Notably, such targeted policy utilizing the current combined 28% capacity of both programs, yields outcomes comparable to policies that accommodate all applicants into either Mentoring or Challenges. This showcases that in this context targeting admission has a very high potential impact.
A noteworthy aspect of Mentoring is that mentors have the autonomy to choose their mentees. To facilitate randomized access for our experiment, we asked mentors to shortlist two candidates and we randomized within these pairs. This resulted in three categories: applicants not chosen by mentors, those shortlisted but placed in the control group through randomization, and those shortlisted and included in the treatment group. Importantly, all applicants underwent an assessment of their technical skills before being presented to mentors. By the end of our observation period, there was a 20 percentage point difference in securing tech jobs between those selected but placed in control and the non-selected applicants, with the latter securing job in technology at the probability of 22%. This underscores that mentors lean towards candidates they deem more promising. While this approach boosts the success rate of Mentoring, reflected in the proportion of participants landing tech jobs, it does not necessarily maximize impact since many of these candidates might have secured tech roles even without Mentoring. Through an off-policy evaluation, we explored potential alternative mentor selection strategies. Our findings suggest that mentors focusing on participants with the greatest treatment effects, rather than those already poised for success, can significantly amplify the efficacy of Mentoring. We estimate that when mentors select 40% of participants based on the highest treatment effect, the share of applicants eventually finding a job in technology is 31% compared to 21% when they choose participants that have the highest outcomes if treated, and 24% when they select participants at random.
@@ Appendix B Covariate balance
In Figure 6 we show the balance of covariates in the mentoring experiment: each dot corresponds to the absolute difference of means in treatment and control divided by the square root of the sum of variances. The only difference that is statistically significant is the graduate degree, where 66% of treated subjects have a graduate degree and 55% of control (p-value 0.05).
@@
In Figure 7 we show the balance of covariates in the Challenges experiment: each dot corresponds to the absolute difference of means in treatment and control divided by the square root of the sum of variances. None of the differences is statistically significant.
@@
Appendix C Short-term average treatment effects estimates from the AIPW estimator
In Table 7, we present the estimates of the average treatment effects using the Augmented Inverse Propensity Weighing estimator (Robins et al., 1994). AIPW is a doubly-robust method which means that it adjusts for covariates in the outcome model and the propensity score. We use the grf implementation of the AIPW estimator, as in Athey et al. (2019). We use the full set of covariates as controls. We find that the estimates from the AIPW estimator are very similar to the estimates obtained using the difference-in-means estimator.
Table 7: Average treatment effect estimates.
Mentoring Challenges
Tech job 4 months New job Tech job 4 months New job
ATE AIPW 0.129 0.037 0.088 0.0437
(0.05) (0.06) (0.04) (0.04)
ATE 0.129 0.045 0.089 0.040
(0.05) (0.05) (0.04) (0.05)
Mean treatment 0.421 0.467 0.284 0.333
(0.04) (0.04) (0.03) (0.03)
Mean control 0.293 0.422 0.196 0.293
(0.04) (0.04) (0.03) (0.03)
No. of obs. treatment 152 152 183 183
No. of obs. control 147 147 225 225
Note: The estimates of the average treatment effects of both the mentoring program and Challenges on the probability of finding a new job in technology and finding any new job. The first row shows the estimates using the augmented inverse propensity weighting (AIPW) estimator. The second row shows estimates of the average treatment effect using the difference-in-means estimator. Rows three and four present mean outcomes in treatment and control groups and rows five and six the number of observations per experimental group. The first and second columns present the results of the mentoring intervention on the probability of finding a job in technology and the probability of finding any new job. Columns three and four show the results of the impact of participation in Challenges on the probability of finding a job in technology and any new job. Standard errors are presented below the average treatment effect estimates. The variance of the difference-in-means estimates the average treatment effect of mentoring estimated using equation 2. The variance of the difference-in-means estimate of the Challenges intervention is estimated using the classical Neyman-Splawa estimator. AIPW variance estimated using grf implementation of AIPW estimator (see Athey et al. (2019) section 4 for details).
Appendix D Long term treatment effects
In this, section we present the estimates from two Cox proportional hazard models for both Mentoring and Challenges. First, we estimate a model with a full set of controls and second, we consider a model where the treatment variable has time-varying coefficients and the full set of controls. Table 8 shows the results. We estimate the coefficient of 0.26 in the case of Mentoring, which corresponds to a 29% increase in the probability of having a tech job. We find higher treatment effects in the case of Challenges, a coefficient of 0.37, corresponding to a 45% increase. We also consider the model in which we analyze the impact of participation in Dare IT programs across time - columns 2 and 4 of Table 8. We confirm the observation from Figure 2, treatment effects (especially in the case of Challenges) are realized predominantly after the end of the two programs. Furthermore, we do not observe attenuation of the treatment effect over time.
Table 8: Long-term treatment effects
Dependent variable: tech job
Mentoring Mentoring time-varying Challenges Challenges time-varying
Treated ${}^{***}$ (0.0501) ${}^{***}$ (0.065)
Treated:during program 0.196 (0.130) --0.074 (0.239)
Treated: 4 months after the program ${}^{***}$ (0.071) ${}^{***}$ (0.094)
Treated: 5 months plus after the program ${}^{***}$ (0.082) ${}^{***}$ (0.096)
Full set of controls x x x x
Observations 4,275 4,275 5,490 5,490
Note: Selected coefficient estimate from the proportional hazard models. The first column is the mentoring experiment. The second column is the mentoring experiment with time-varying coefficients. The third column is the challenges experiment. In the last column, the challenges experiment with time-varying coefficients. In Appendix G, we present the Cox proportional hazard model diagnostics.
Appendix E Comparison of salary levels from Glassdoor
@@
Appendix F Outcomes measured in surveys
@@
Appendix G Proportional hazards assumption
@@
Appendix H Bounds on the average treatment effect
In this section, we perform three robustness checks on the average treatment effects. First, we consider the AT bounds method introduced by Lee and Weidner (2021). Second, due to changes in LinkedIn profile URLs we could not collect outcome data from 17 Challenges participants. To obtain bounds on the treatment effects, we consider two extreme cases - none of the 17 subjects found a job in technology, and all of them did. We estimate treatment effects in both cases. Lastly, we had four participants drop out of the Mentoring program because they secured jobs in the tech industry before the start of the program. We establish bounds based on this information, assuming two distinct scenarios: (i) all corresponding control group participants also found jobs in technology, and (ii) none of them did.
Lee and Weidner (2021) proposes bounds on the average treatment effects that are designed to be robust in situations where there are many conditioning variables, or they have many values, or when the overlap condition is violated. We use R implementation of their algorithm (Lee and Weidner, ).
In the case of Mentoring we consider the following variables: Social science, STEM, Family or Friends in IT, Mother, Applied Before, First-time mentor, Managerial Experience, Career changer, Mentor’s Years of Experience, Professional Experience. We estimate the bounds on the average treatment effect to be 0.11 to 0.18 with a confidence interval of 0.03 to 0.26. In the case of Challenges, we consider the following characteristics: social science, STEM, Specialization, Residence, Age, Professional experience, Degree. We obtain bounds of 0.10 to 0.12 with a confidence interval of 0.02 to 0.20.
Next, we address the unavailability of LinkedIn data for 17 Challenges participants. We explore two extreme scenarios: one where all 17 did not secure tech-sector jobs and another where all 17 did. We estimate the average treatment effect in these scenarios using a difference-in-means estimator, without controlling for observable features due to the lack of LinkedIn data. When we assign zero outcomes for the missing subjects, the estimated average treatment effect is 6.6 percentage points, with a confidence interval ranging from -1 to 15 percentage points. Assuming positive outcomes for all missing variables, we estimate the average treatment effect at 15 percentage points, with a confidence interval spanning 7 to 20 percentage points.
Finally, we conduct a robustness check of our evaluation of the Mentoring experiment by examining two scenarios: (i) all four control group subjects also found tech jobs, and (ii) none of them did. We estimate treatment effects using the difference-in-means estimator. In the pessimistic scenario, the estimated treatment effect is 12 percentage points, with a confidence interval of 2 to 23 percentage points. In the optimistic scenario, we estimate the treatment effect at 15 percentage points, with a confidence interval spanning 4 to 26 percentage points.
Appendix I Heterogeneous treatment effects
In Table 10 we show estimates from the Cox model including the full list of interactions.
Table 10: Heterogeneous treatment effects of the interventions.
Dependent variable: tech job

### v4

@@ 6.3 Counterfactual Mentors’ Selection Policies
We use our off-policy tools to examine alternative selection strategies that mentors could adopt and their potential implications for the mentoring program’s effectiveness. Specifically, we consider four scenarios: (i) Highest CATE—targeting applicants predicted to have the highest treatment effects, (ii) Most Promising—targeting applicants expected to have the highest outcomes without participating in Mentoring, (iii) Least Promising—selecting applicants expected to have the lowest outcomes if they do not participate, and (iv) Highest Chances if Treated—selecting participants with the highest predicted outcomes if they participate in Mentoring. Additionally, we include two benchmarks: mentors selecting at random (Random Choice) and the observed allocation (Observed). In each of the four policies, we assume that 15% of applicants participate in Mentoring, which is an approximate share of participants to applicants observed by DareIT in recent years.
Table 9 examines the extent to which different allocation policies select the same individuals Each cell shows the percentage of individuals in the row policy who also appear in the column policy. First, we consider the relationship between mentor-based selection (Current) and the Highest CATE algorithmic policy. Among current mentor-selected participants, 70.4% would also be assigned to Mentoring under the Highest CATE algorithm, demonstrating substantial concordance between expert judgment and data-driven targeting (note that we do not use the information on whether a candidate was selected or not in the policy estimation).
Current mentor selections align most closely with Highest CATE (70.4%), followed by Most Promising (29.6%), with little overlap with Least Promising (18.3%) or Highest Chances (19.7%). This pattern suggests mentors implicitly optimize for treatment effects rather than baseline characteristics or final outcomes, though with some cream-skimming tendency. The 29.6% overlap with Most Promising indicates approximately one-third of mentor selections favor strong baseline candidates, while the predominant 70% overlap with Highest CATE confirms that mentors primarily target based on perceived program responsiveness. The low overlap with Least Promising (18.3%) suggests mentors, whether by design or implicit bias, rarely select individuals with the weakest baseline labor market attachment, even when such individuals might benefit from treatment.
Table 9: Overlap Between Allocation Policies
Policy H. CATE Most Prom. Least Prom. H. Chances Current
Highest CATE 100.0
Most Promising 28.5 100.0
Least Promising 18.2 0.0 100.0
Highest Chances 27.9 26.7 31.5 100.0
Current 70.4 29.6 18.3 19.7 100.0
Notes: Each cell shows the percentage of individuals in the row policy that also appear in the column policy. Only lower triangle shown to avoid redundancy; upper triangle omitted. Sample sizes in parentheses after policy names. Diagonal entries are 100% by construction (bold). H. CATE = Highest CATE; H. Chances = Highest Chances; Most/Least Prom. = Most/Least Promising.
Table 10 compares treatment effects across six allocation policies for the Mentoring program. The results demonstrate that targeting based on predicted treatment effects substantially outperforms alternative strategies. The Highest CATE policy achieves treatment effects of 3.59-3.59 months (SE = 0.033), the largest magnitude across all policies. Random allocation serves as the baseline benchmark, generating treatment effects of 2.20-2.20 months (SE = 0.040). Current mentor-based selection performs remarkably well, achieving 3.25-3.25 months (SE = 0.100), a 48% improvement over random allocation. This validates that mentors successfully identify high-benefit participants through their evaluative selection process.
Table 10: Allocation Policy Comparison: Mentoring Program
Policy Baseline Time Treatment Effect
(months) (months)
Current 11.09 (0.131) -3.25 (0.100)
Random 11.36 (0.059) -2.20 (0.040)
Data-driven policies:
Highest CATE 11.08 (0.086) -3.59 (0.033)
Most Promising 9.71 (0.044) -2.27 (0.075)
Least Promising 13.44 (0.085) -1.94 (0.069)
Highest Chances 11.77 (0.152) -2.14 (0.079)
Notes: Comparison of allocation policies for the Mentoring program. Baseline Time is the predicted months to employment without treatment (lower is better); Treatment Effect is the predicted change in months from treatment (more negative = faster employment). Standard errors in parentheses. Current reflects mentor-selected participants. Random is random assignment benchmark. Highest CATE maximizes treatment effects; Most/Least Promising target best/worst baseline outcomes; Highest Chances minimizes post-treatment time to employment.
Notably, the Highest CATE algorithmic policy achieves an additional 10.4% improvement over mentor selection, increasing treatment effects from 3.25-3.25 to 3.59-3.59 months. We obtain the 0.34-month gain using only six readily-observable characteristics (education, location, age, experience) while mentors access rich soft information from interviews and holistic assessment. The overlap analysis (Table 9) reveals that 70.4% of mentor-selected participants would also be assigned under Highest CATE, demonstrating substantial concordance between expert judgment and data-driven targeting. The finding that simple observable characteristics can meaningfully improve upon expert selection—even when experts demonstrate strong performance—underscores the value of systematic data-driven approaches as complements to human judgment.
Policies optimizing objectives other than treatment effects yield substantially weaker results. Most Promising, which targets applicants with the strongest baseline prospects (9.71 months to employment), achieves treatment effects of only 2.27-2.27 months, barely exceeding random allocation despite successfully cream-skimming the highest-potential candidates. This demonstrates that strong baseline prospects do not predict strong treatment responsiveness. Conversely, Least Promising targets individuals with weakest baselines (13.44 months) but generates the smallest treatment effects (1.94-1.94 months), revealing an unfortunate equity-efficiency tradeoff: those who need help most benefit least from the program. Highest Chances, which optimizes final employment outcomes, achieves 2.14-2.14 months, marginally worse than random assignment. These comparisons underscore that maximizing program impact requires explicitly targeting predicted treatment effects rather than baseline characteristics or final outcomes, as treatment response heterogeneity does not align with baseline heterogeneity.
@@ 7 Conclusion
We show that a mentoring program and an online training program called Challenges aimed at developing labor market signals increase the probability of having a job in the technology sector by over 30% twelve months after the end of the program. We estimate that Mentoring increases technology employment from 46% to 61%, while Challenges increases it from 27% to 38%, with treatment effects persisting over the full observation period. Furthermore, the online training program developed during this research project is cheap to operate, at approximately \$15 per person, and suitable for scaling up, suggesting that it can have an impact extending beyond the current setting.
@@
Our findings reveal strong patterns of heterogeneity in treatment effects. The benefits derived from the two programs depend on a number of participants’ characteristics including the educational background, length of professional experience, type of residence, and age. For example, we find that older participants that come from small towns have particularly high treatment effects in both programs, especially in Mentoring. Furthermore, certain groups benefit distinctly from one program over the other. Specifically, participants with short professional experience benefit from Challenges, while those with over five years of experience from Mentoring. These differential patterns provide evidence on the underlying mechanisms: Mentoring expands professional networks (increasing LinkedIn connections by 52.7, a 23% increase), while Challenges provides credible skill signals. Notably, network expansion occurs only in Mentoring, and heterogeneity patterns align with each program’s mechanism: those with weaker networks benefit most from Mentoring, while those with weaker credentials benefit most from Challenges.
@@
For the foreseeable future, the demand for the two programs, particularly for Mentoring will exceed the number of available spots, and Dare IT will grapple with the question of whom to admit. We propose an admission policy that maximizes Dare IT’s impact, by prioritizing applicants with the highest predicted treatment effects. Consistent with the high degree of heterogeneity in the effectiveness of the two programs we find that such targeted admission policies can substantially increase Dare IT’s effectiveness. Specifically, we estimate that optimal targeting achieves outcomes 86% larger than random assignment at current capacity levels (30% of applicants). Notably, the value of targeting persists even as capacity expands: at 50% capacity per program, targeted allocation maintains outcomes 23% larger than random assignment. This demonstrates that targeting generates value both through selecting who participates and through matching individuals to their highest-benefit program. The complementarity between individual characteristics and program type creates allocation gains beyond simple capacity expansion.
A noteworthy aspect of Mentoring is that mentors have the autonomy to choose their mentees. To facilitate randomized access for our experiment, we asked mentors to shortlist two candidates and we randomized within these pairs. Comparing selected-but-control applicants to non-selected applicants reveals that mentors successfully identify promising candidates: mentor-selected participants achieve outcomes 48% larger than random allocation. However, mentors appear to prioritize candidates with strong baseline prospects, those likely to secure technology employment even without the program, rather than those who would benefit most from participation. We evaluate an alternative algorithmic allocation policy that selects participants based on predicted treatment effects using only six readily-observable characteristics (education, location, age, experience). Despite mentors’ rich soft information from interviews and holistic assessment, algorithmic targeting achieves outcomes 63% larger than random allocation, representing an additional 11% improvement over mentor judgment. This finding is particularly striking given that algorithmic and mentor-based selections exhibit 70% concordance: the algorithm largely validates expert choices while identifying additional high-benefit candidates. The results demonstrate that systematic patterns in treatment response exist even after expert selection captures most available heterogeneity, and that simple demographic targeting can complement human judgment to enhance program effectiveness.
@@ Appendix B Covariate balance
In Figure 5 we show the balance of covariates in the mentoring experiment: each dot corresponds to the absolute difference of means in treatment and control divided by the square root of the sum of variances.
@@
In Figure 6 we show the balance of covariates in the Challenges experiment: each dot corresponds to the absolute difference of means in treatment and control divided by the square root of the sum of variances. The only variable that has a statistically significant differnece is social science.
@@
Appendix C Kaplan Meier Survival Curves
We present Kaplan Meier estimates of the survival function $S(t)$, defined as the probability of not yet obtaining a technology job by month $t$.
Figure 7 shows the survival functions for the Mentoring program. Relative to both the control group and applicants, participants in the Mentoring treatment group find technology jobs more rapidly: their survival curve falls more steeply, indicating earlier entry into technology jobs. The difference between treated and control groups emerges within the first few months and grows steadily over the 15-month horizon. We can also notice that there is a substantial gap between the applicants and both experimental groups which emerges already during the first four months,
Figure 7: Survival Curves Mentoring Program.
Note: Kaplan Meier survival curves for the Mentoring program, including treated participants, control group, and applicants. A steeper decline indicates earlier job placement.
Figure 8 presents analogous estimates for the Challenges program. Treated participants also transition into jobs more quickly than the control group or applicants, though the gap is smaller in magnitude than in the Mentoring program.
Figure 8: Survival Curves Challenges Program.
Note: Kaplan Meier survival curves for the Challenges program, including treated participants, control group, and applicants. Treated participants transition to jobs earlier, though the effect is smaller than in Mentoring.
Appendix D Comparison of salary levels from Glassdoor
@@
Appendix E Outcomes measured in surveys
@@
Appendix F Proportional hazards assumption
@@
Appendix G Bounds on the Average Treatment Effect Estimates
In this section, we perform robustness checks on the average treatment effect estimates. First, we consider several cases in which we change which subjects are included in the analysis. Second, we estimate AT bounds following lee2021bounding. Third, we estimate the average treatment effects at 4 months after the programs.
Robustness to Sample Construction and Attrition.
Our main analysis excludes participants who secured technology employment between randomization and program launch, as well as mentors who declined to participate in the experimental design. While these exclusions yield conservative estimates, we verify that our conclusions are robust to alternative sample definitions. Table 12 presents treatment effect estimates under seven alternative specifications. For each, we report sample sizes, mean outcomes by treatment status, and the estimated average treatment effect using difference-in-means without covariate adjustment. All estimates use the twelve-month tech employment outcome.
We consider four specifications for Mentoring. First, our baseline estimate (15.3 percentage points, SE 5.7) uses 151 treatment and 147 control participants, reflecting the 4 mentor pairs where both candidates were assigned to treatment due to control participants securing early employment. Second, restricting to the 126 perfectly matched pairs yields a slightly smaller but still substantial effect (11.9 percentage points, SE 6.3). Third, including the 4 participants who found jobs before program launch—coding their outcomes as 1—increases the estimated effect to 16.4 percentage points. Fourth, additionally including the 9 non-participating mentors under the assumption their hypothetical mentees would not have found jobs reduces the estimate to 13.0 percentage points. All four specifications yield treatment effects between 11.9 and 16.4 percentage points, each statistically significant at the 5% level.
For Challenges, we examine three specifications. Our baseline estimate (11.3 percentage points, SE 4.8) excludes 17 treatment participants who secured technology jobs before program launch. Including these individuals and coding their outcomes as 1 increases the effect to 17.1 percentage points—as expected, since we add successful job-finders to the treatment group. Finally, we address the 17 treatment and 7 control participants with missing twelve-month LinkedIn data. Under the conservative assumption that all missing treatment observations had outcomes of 0 and all missing control observations had outcomes of 1—the scenario most unfavorable to finding a treatment effect—we still estimate a positive effect of 10.5 percentage points (SE 4.6, p=0.02).
Table 12: Robustness to Sample Construction and Attrition
Specification Sample Size Mean Outcome ATE
Treated Control Treated Control (SE)
Panel A: Mentoring
   Main specification 151 147 0.609 0.456 0.153***
   Balanced pairs only 126 126 0.603 0.484 0.119*
   Include pre-program employment 155 147 0.619 0.456 0.164***
   Include non-participants 164 147 0.585 0.456 0.130**
Panel B: Challenges
   Main specification 166 218 0.380 0.266 0.113**
   Include pre-program employment 183 218 0.437 0.266 0.171***
   Pessimistic attrition bounds 200 227 0.400 0.295 0.105**
Note:Estimates of average treatment effects on twelve-month technology employment under alternative sample definitions. Main specifications correspond to Table 2. Balanced pairs only (Mentoring) restricts to the 126 mentor pairs with both treatment and control observations. Include pre-program employment adds participants who secured technology jobs between randomization and program launch. Include non-participants (Mentoring) adds 9 mentors who declined experimental participation, assigning their hypothetical mentees outcomes of 0. Pessimistic attrition bounds assign outcomes of 0 to 17 missing treatment observations and 1 to 7 missing control observations. All estimates use difference-in-means without covariate adjustment. Standard errors in parentheses. *** p<0.01, ** p<0.05, * p<0.1.
These robustness checks establish three key points. First, our main estimates are insensitive to reasonable alternative sample construction decisions, with all specifications yielding economically large and statistically significant treatment effects. Second, the range of point estimates across specifications (11.9 to 16.4 for Mentoring, 10.5 to 17.1 for Challenges) provides additional bounds on the true treatment effects that are consistent with our main findings. Third, even under extremely conservative assumptions about missing data—assumptions, the Challenges program shows significant positive impacts. This addresses potential concerns that selective attrition might bias our twelve-month estimates.
Lee Bounds for Selective Attrition.
Our twelve-month outcome measurement experienced modest attrition in the Challenges program: 6% had missing LinkedIn data due to profile URL changes or privacy setting modifications. We assess whether selective attrition could bias our estimated treatment effect using the bounding procedure of lee2009training. We consider the most conservative case, in which subjects that found jobs before the start of the program are not considered; thus, we focus only on the attrition during the program, due to missing LinkedIn data.
The Lee bounding procedure addresses selective attrition by considering worst-case scenarios about which individuals are missing. Since treatment experienced higher attrition, the conservative assumption for estimating positive effects is that the best-performing treatment participants were most likely to attrite (e.g., successful job-finders who deactivated LinkedIn). To construct the lower bound, we trim 6.8% of observations from the upper tail of the observed treatment outcome distribution, the proportion representing excess attrition relative to control, and compare the mean of this trimmed distribution to the full control group mean. The upper bound uses all observed treatment outcomes compared to all observed control outcomes. We estimate Lee bounds of [0.065, 0.114]. The upper bound of 0.114 is nearly identical to our main point estimate of 0.113, while the positive lower bound of 0.065 indicates that even under worst-case assumptions about selective attrition, the Challenges program effect remains economically meaningful. The 95% confidence interval on the lower bound is [-0.029, 0.159]; while this marginally includes zero due to sampling variability in the observed outcomes, the point estimate is bounded away from zero.
Average Treatment Effect Estimates at 4 Months.
Our main analysis uses employment outcomes measured twelve months post-program. However, we were unable to collect LinkedIn data for 17 Challenges treatment participants and 7 control participants at the twelve-month measurement due to profile URL changes or privacy setting modifications. A natural concern is whether selective attrition biases our estimated treatment effects. To address this concern, we re-estimate average treatment effects using employment outcomes measured four months after program completion. At this earlier measurement point, we successfully collected complete LinkedIn data for all randomized participants in both programs. Any attrition-induced bias is therefore eliminated at this time horizon.
Table 13 presents the results. We find substantively similar treatment effects at the four-month horizon compared to our main twelve-month estimates (reported in Table 2). For Mentoring, the four-month treatment effect is 15.3 percentage points (SE 5.7), nearly identical to the twelve-month effect of 14.5 percentage points. For Challenges, the four-month effect is 8.9 percentage points (SE 4.3), compared to 11.2 percentage points at twelve months. The larger point estimate at twelve months suggests that the Challenges program’s impact may strengthen over time as participants leverage their portfolios in job applications.
Table 13: Average Treatment Effects of Mentoring and Challenges at 4 Months.